Your friend, Suzie Whitson, has designed a new type of outdoor toy that helps children learn basic concepts such as colors, numbers, and shapes. Suzie’s product will target two groups: day care centers in warm climates and home school programs. Her company is Jiffy Jet and costs for last month follow:

Factory rent $ 3,070
Company advertising 1,110
Wages paid to assembly workers 31,600
Depreciation for salespersons’ vehicles 2,020
Screws 585
Utilities for factory 840
Assembly supervisor’s salary 3,570
Sandpaper 135
President’s salary 5,070
Plastic tubing 4,100
Paint 245
Sales commissions 1,270
Factory insurance 1,050
Depreciation on cutting machines 2,040
Wages paid to painters 8,000

Determine the total cost for each of the following:

a. Direct Materials
b. Direct Labor
c. Manufacturing Overhead
d. Prime Cost
e. Conversion Cost
f. Total Product Cost
g. Total Period Cost
h. Total Variable Cost
i. Total Fixed Cost
Answer #1

a. Direct material = Plastic tubing

Direct material = $4,100

b. Direct labor = Wages paid to assembly workers + Wages paid to painters

Direct labor = $31,600 + 8,000 = $39,600

c. Manufacturing overhead = Factory rent+Screws+Utilities for factory+Assembly supervisor salary+Sandpaper+Paint+Factory insurance+Depreciation on cutting machines

Manufacturing overhead = $3,070+585+840+3,570+135+245+1,050+2,040 = $11,535

d. Prime cost = Direct material + Direct labor

Prime cost = $4,100 + 39,600 = $43,700

e. Conversion cost = Direct labor + Manufacturing overhead

Conversion cost = $39,600 + 11,535 = $51,135

f. Total product cost = Prime cost + Manufacturing overhead

Total product cost = $43,700 + 11,535 = $55,235

g. Total period cost = Company advertising+Depreciation for salesperson vehicle+Presidents salary+Sales commission

Total period cost = $1,110+2,020+5,070+1,270 = $9,470

h. Total variable cost = Wages paid to assembly workers+Srews+Utilities for factory+Sandpaper+ Plastic tubing+Paint+Sales commission+Wages paid to painters

Total variable cost = $31,600+585+840+135+4,100+245+1,270+8,000 = $46,775

i. Total fixed cost = Factory rent+Company advertising+Depreciation for salesperson vehicle+Assembly supervisor salary+Presidents salary+Factory insurance+Depreciation on cutting machines

Total fixed cost = $3,070+1,110+2,020+3,570+5,070+1,050+2,040 = $17,930